# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v2

@@ Abstract
Recent methodological developments have introduced new black-box approaches to better estimate heterogeneous treatment effects; however, these methods fall short of providing interpretable characterizations of the underlying individuals who may be most at risk or benefit most from receiving the treatment, thereby limiting their practical utility. In this work, we introduce causal distillation trees (CDT) to estimate interpretable subgroups. CDT allows researchers to fit any machine learning model to estimate the individual-level treatment effect, and then leverages a simple, second-stage tree-based model to “distill” the estimated treatment effect into meaningful subgroups. As a result, CDT inherits the improvements in predictive performance from black-box machine learning models while preserving the interpretability of a simple decision tree. We derive theoretical guarantees for the consistency of the estimated subgroups using CDT, and introduce stability-driven diagnostics for researchers to evaluate the quality of the estimated subgroups. We illustrate our proposed method on a randomized controlled trial of antiretroviral treatment for HIV from the AIDS Clinical Trials Group Study 175 and show that CDT out-performs state-of-the-art approaches in constructing stable, clinically relevant subgroups.
11footnotetext: The authors would like to thank Sam Pimentel, Yan Shuo Tan, and participants at the American Causal Inference Conference, Johns Hopkins Biostatistics Causal Inference Seminar, UNC Chapel Hill Biostatistics Seminar for their feedback and comments.22footnotetext: Denotes equal contribution.33footnotetext: Yale University. Email: melody.huang@yale.edu, URL: www.melodyyhuang.com44footnotetext: University of Notre Dame. Email: ttang4@nd.edu, URL: https://tiffanymtang.github.io/55footnotetext: University of California Irvine. Email: anamaria.kenney@uci.edu
@@ 1 Introduction
In its simplest form, many subgroup analyses in causal inference focus on pre-specified subgroups based upon prior substantive knowledge. However, it can still be challenging to reason about higher-order interactions between covariates. In the absence of prior knowledge, different tree-based approaches have been proposed to discover new subgroups with heterogeneous treatment effects (e.g., Su et al., 2009; Lipkovich et al., 2011; Loh et al., 2015), of which causal trees (Athey & Imbens, 2016) are arguably the most popular. [^1] Citation count was used as a proxy for popularity. In comparing the citation count of the different papers, as of 2025, the original causal tree paper by Athey & Imbens (2016) has over 2,000 citations, while the alternative methods have anywhere from 100-500 citations. Similar to decision trees (Breiman et al., 1984), causal trees partition the underlying covariate space into groups by finding clusters of individuals who have similar treatment effects. Moreover, unlike linear regression approaches, causal trees output a readily-interpretable partition of subgroups while also accounting for potential high-order interactions.
@@
In the following paper, we propose causal distillation trees (CDT) to stably estimate interpretable subgroups. CDT allows researchers to leverage the power of black-box meta learners, while preserving the interpretability of a simple tree output. CDT is a two-stage learner. The first stage estimates a heterogeneous treatment effect model using a flexible metalearner to generate a predicted individual-level treatment effect for each unit. The second stage then distills the information from the predicted individual-level treatment effect through a decision tree to construct interpretable subgroups. The first step serves as a de-noising step to help mitigate the instabilities that decision trees traditionally suffer from. The second step of ‘distillation’ draws upon existing ideas in machine learning (e.g., Hinton, 2015; Menon et al., 2021; Dao et al., 2021). However, unlike typical machine learning settings, where the primary goal is to improve the second stage learner’s predictive ability, we leverage distillation to improve our ability to estimate interpretable subgroups using the second stage decision tree learner.
@@ 2 Setup and Notation
[…]
@@
$\tau_{\textsc{ATE}}=\frac{1}{n}\sum_{i=1}^{n}\left\{Y_{i}(1)-Y_{i}(0)\right\}.$
@@
where $\Omega$ represents a class of partitions in the covariate space of $X$. Under squared loss, the optimal partition corresponds to a subset in the covariate space of $X$ that maximizes the difference in treatment effect heterogeneity across subgroups. In other words, within a partitioned group, the treatment effect heterogeneity is minimized, thereby maximizing the variance in treatment effects across groups. In practice, researchers can choose alternative loss functions that correspond to what are substantively meaningful partitions in the covariate space. However, the theoretical guarantees we derive utilize the squared loss, and generalizing the results for other loss functions is an open avenue of future research.
@@ 3.1 Overview
We propose a method, causal distillation trees (CDT), which stably estimates interpretable subgroups. CDT is a two-stage learner (Figure 1 and Algorithm 1). The first stage learner, referred to as the teacher model, learns an informative model for the individual-level treatment effect using the observed data $\{Y,\;Z,\;X\}$. Using this teacher model, we generate a prediction $\hat{\tau}_{i}^{d}$ for the individual-level treatment effect for all units. In the second stage, we distill the predicted $\hat{\tau}_{i}^{d}$’s into interpretable subgroups by fitting a decision tree, called the student model, to predict the $\hat{\tau}_{i}^{d}$’s from the covariates $X$. The estimated decision rules from the decision tree define a discrete partition, which naturally maps to the collection of subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\;\ldots,\;\hat{\mathcal{G}}_{G}(X)\right\}$ (see Figure A1). Finally, using the estimated partitions, we estimate the subgroup ATE for each subgroup.
Figure 1: Overview of Causal Distillation Trees (CDT). CDT leverages a two-stage procedure, which first fits a teacher model to estimate individual-level treatment effects and then fits a student model (e.g., a decision tree) to distill the estimated individual-level treatment effects and produce interpretable subgroups. The two-stage learner is learned using the training data (blue-gray boxes). Using the estimated subgroups, the subgroup ATEs are honestly estimated with a held-out estimation set (yellow box).
Informally, the first stage learner smooths the individual-level treatment effects by projecting $\tau_{i}$ into the basis space of the covariates $X$. By using the projected version of the individual-level treatment effect instead of $\tau_{i}$, the second stage learner will be able to more stably estimate the subgroups, as the teacher model will have de-noised the outcomes.
@@ 3.2 Consistency of subgroup estimation: Single rule setting
Let $F_{X}$ be the cumulative distribution function of the covariate $X$. Define the points $s_{l}$ and $s_{u}$ such that $10<F_{X}(s_{l})<F_{X}(s_{u})<$ 1. Let $s$ be an optimal split as defined in Definition 3.1. Assume that (i) $F_{X}$ is absolutely continuous with a density $dF_{X}$ that is bounded away from zero and is continuous in a neighborhood of $s$, and (ii) $s$ is unique.
@@
Under this validity of the teacher model, the population-level optimal subgroups remain unchanged after distilling the individual-level treatment effects. In other words, the optimal population splits from the predicted individual-level treatment effects $\hat{\tau}_{i}^{d}$ are equivalent to the optimal split with the true individual-level treatment effects $\tau_{i}$. A sufficient, though not necessary, condition for a valid teacher model is a correctly-specified teacher model, such that the residual noise that cannot be explained by the model must be uncorrelated with the covariates $X$ (i.e., exogeneity). In randomized control trials, we expect this to be true if the teacher model is flexible enough and the sample size is sufficiently large. Assumption (3.2) can also be relaxed to consider approximately valid teacher models (see Appendix A.2 for additional discussion).
@@
  • $\tau(x)$ is continuous in $(l,s)$ and $(s,u)$, where $\lim_{x\to s^{-}}\tau(x)=\tau(s^{-})\text{ and }\lim_{x\to s^{+}}\tau(x)=\tau(s^{+})$.
  • For $\epsilon>0$, there exists an $21<\alpha<$ 2 where: $\inf_{|s^{\prime}-s|<\epsilon,s\neq s^{\prime}}\frac{|H(s)-H(s^{\prime})|}{|s-s^{\prime}|^{\alpha}}>0$.
@@
Proposition 3.1 formalizes that the rate of convergence depends directly on how smooth the underlying data generating process is. The rate of convergence in the estimated subgroups inherits a characteristic highlighted in Escanciano (2020), which bridged the gap in convergence rate of the estimated split in a decision tree $\hat{s}_{n}$ between a fully continuous setting (considered in (Banerjee & McKeague, 2007; Buhlmann & Yu, 2002)) and discontinuous settings (considered in (Chan, 1993; Kosorok, 2008)).
@@
In addition to consistently recovering the subgroups, CDT can further improve the stability of the split estimation by de-noising the individual-level treatment effect and increasing the signal strength in the individual-level treatment effect. To illustrate, in the following example, we consider a setting in which researchers have access to the true individual-level treatment effect $\tau_{i}$ and estimate a decision tree using the covariates $X_{i}$. In practice, this is infeasible, as researchers can never observe $\tau_{i}$, but this comparison serves as a helpful benchmark to compare the potential improvements from distillation.
@@ 3.3 General setting with multiple rules
For each subgroup index $g\in\{1,\ldots,G\}$ and rule index $r\in\{1,\ldots,r_{g}\}$, without loss of generality, write the subgroup defined by the $r-1$ previously selected binary decision rules as $\mathcal{G}_{g}^{(r-1)}(X)=\prod_{k=1}^{r-1}\mathbbm{1}\left\{X^{(j_{k})}\lesseqgtr s^{(j_{k})}\right\}$. We assume that there is separability, conditional on the previously selected binary decision rules:
@@
Assumption (3.5) ensures that the underlying subgroups have an optimal substructure so that a greedy second-stage decision tree learner is valid. If Assumption (3.5) is not met, standard greedy tree algorithms like CART can fail to recover the optimal partitions even with infinite amounts of data (see (Tan et al., 2024) for a formal discussion). Recent work has introduced alternative algorithms to construct optimal trees under weaker assumptions, but at a much higher computational cost, and are often restricted to specific settings such as binary features, with limited out-of-sample performance gains relative to greedy trees (e.g., Hu et al., 2019; van der Linden et al., 2024). Given the popularity of greedy search approaches and the computational burden of many optimal tree construction approaches, we focus our discussion on settings in which greedy search algorithms can feasibly construct optimal subgroups.
@@
$\mathbb{E}\left(\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert\right)\leq$ $\frac{2r_{g}(C_{\tau}+M\delta)}{\delta}\left\lvert\hat{s}^{(k)}_{n}-s^{(k)}\right\rvert+$
@@
The results of Theorem 3.1 follow from first noting that the estimated splits will converge in probability to the optimal split (see Lemma 3.1), and then showing that the within-sample loss will converge in probability to the true population-level loss (see Appendix B) In particular, the rate of convergence of $\hat{\mathcal{G}}_{g}(X_{i})$ to $\mathcal{G}_{g}(X_{i})$ is dependent on (1) the degree of separability in the losses between relevant and irrelevant rules (i.e., $\delta$), and (2) the underlying convergence rate of the sample splits to the true optimal splits (i.e., the rate at which $\left\lvert\hat{s}_{n}^{(k)}-s^{(k)}\right\rvert\to 0$). An immediate implication of Theorem 3.1 is that not only will the tree consistently split at the correct cutoff points, but it will also correctly select the important features that should be used in the splitting criteria.
@@
Figure 2: Comparing the performance of causal trees versus CDT (using causal forests as the teacher model), measured via (A) number of estimated subgroups as well as (B) number of true positive and (C) number of false positive features used in the estimated subgroups. The oracle number of subgroups, true positives, and false positives are shown as dashed gray lines. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@ 3.4 Estimating subgroup average treatment effects
Under Assumptions (3.1)-(3.7), the subgroup difference-in-means estimator will be an unbiased estimator for the sample subgroup ATE. Moreover, the variance of $\hat{\tau}^{(g)}$ is
$\text{var}(\hat{\tau}^{(g)})=\frac{1}{n_{g}}\left\{\beta_{g}(1)\text{var}_{g}(Y_{i}(1))+\beta_{g}(0)\text{var}_{g}(Y_{i}(0))-\text{var}_{g}(\tau_{i})\right\},$
where $\beta_{g}(z)=\mathbb{E}\{1/\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\}$ and $\text{var}_{g}\left\{Y(z)\right\}:=\mathbb{E}\left\{\frac{1}{n_{g}-1}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})(Y_{i}(z)-\right.$ $\left.\overline{Y_{g}(z)})^{2}\right\}$ for $z\in\{0,1\}$.
If researchers are interested in the population-level subgroup ATE, then there will be some finite-sample bias that arises from slippage between the estimated subgroups $\hat{\mathcal{G}}(X)$ and the true optimal subgroups $\mathcal{G}(X)$; however, the subgroup difference-in-means estimator will still be a consistent estimator for the population-level subgroup ATE. (See Appendix A.4 for more discussion.)
From Theorem 3.2, we can also construct a conservative variance estimator using the sample analogs for $\text{var}_{g}(Y_{i}(z))$ for $z\in\{0,1\}$:
@@
where $\hat{\beta}_{g}(z)=\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\right)^{-1}$. Miratrix et al. (2013) showed that the gap between $\hat{\beta}_{g}(z)$ and $\beta_{g}(z)$ can be upper bounded as a function of $1/n$. Asymptotically, this will be equivalent to the robust Huber–White variance estimator.
To consider the asymptotic distribution of the subgroup difference-in-means estimator, we must account for the fact that as the sample size increases, the underlying subgroup allocation will also change. We leverage recent work in developing central limit theorem results for finite populations (e.g., Schochet, 2024; Li & Ding, 2017) to show that under mild regularity conditions, the subgroup difference-in-means estimator will be asymptotically normal.
Theorem 3.3 (Asymptotic Normality of Subgroup Difference-in-Means Estimator).
Assume that as the sample size $n\to\infty$, the total number of subgroups $G$ remain fixed. Furthermore, assume the proportion of units in each subgroup converges to a proportion $\pi_{g}^{*}\in(0,1)$ (i.e., $n_{g}/n\to\pi_{g}^{*}$, where $10<\pi_{g}^{*}<$ 1 for all $g\in\{1,\ldots,G\}$ and $\sum_{g=1}^{G}\pi_{g}^{*}=1$), and the proportion of treated units converges to $p_{z}$, where $p_{z}\in(0,1)$ (i.e., $\frac{1}{n}\sum_{i=1}^{n}Z_{i}\to p_{z}$). Then, for $z\in\{0,1\}$ and $g\in\{1,\ldots,G\}$, if the following condition holds:
The results of Theorem 3.3 can be viewed as a special case of Schochet (2024), Theorem 1. The condition in (5) is a Lindeberg-type condition, which restricts the tail behavior of the underlying potential outcome distribution and is relatively weak.
Following Theorem 3.3, we can apply standard nonparametric tests of treatment effect heterogeneity (see Appendix A). If researchers are interested in evaluating whether the largest treatment effect across the subgroups is statistically significantly different than the other treatment effects, we caution that a different test is needed due to post-selection bias. See (Wei et al., 2024) for more discussion on valid re-sampling procedures for inference in these settings.

### v3

@@ Abstract
Recent methodological developments have introduced new black-box approaches to better estimate heterogeneous treatment effects; however, these methods fall short of providing interpretable characterizations of the underlying individuals who may be most at risk or benefit most from receiving the treatment, thereby limiting their practical utility. In this work, we introduce causal distillation trees (CDT) to estimate interpretable subgroups. CDT allows researchers to fit any machine learning model to estimate the heterogeneous treatment effect, and then leverages a simple, second-stage tree-based model to “distill” the estimated treatment effect into meaningful subgroups. As a result, CDT inherits the improvements in predictive performance from black-box machine learning models while preserving the interpretability of a simple decision tree. We derive theoretical guarantees for the consistency of the estimated subgroups using CDT, and introduce stability-driven diagnostics for researchers to evaluate the quality of the estimated subgroups. We illustrate our proposed method on a randomized controlled trial of antiretroviral treatment for HIV from the AIDS Clinical Trials Group Study 175 and show that CDT out-performs state-of-the-art approaches in constructing stable, clinically relevant subgroups.
11footnotetext: The authors would like to thank Naoki Egami, Erin Hartman, Kosuke Imai, Sam Pimentel, Yan Shuo Tan, and participants at the American Causal Inference Conference, Johns Hopkins Biostatistics Causal Inference Seminar, UNC Chapel Hill Biostatistics Seminar for their feedback and comments.22footnotetext: Denotes equal contribution.33footnotetext: Yale University. Email: melody.huang@yale.edu, URL: www.melodyyhuang.com44footnotetext: University of Notre Dame. Email: ttang4@nd.edu, URL: https://tiffanymtang.github.io/55footnotetext: University of California Irvine. Email: anamaria.kenney@uci.edu
@@ 1 Introduction
In its simplest form, many such subgroup analyses in causal inference focus on pre-specified subgroups based upon prior substantive knowledge. However, it can still be challenging to reason about higher-order interactions between covariates. In the absence of prior knowledge, different tree-based approaches have been proposed to discover new subgroups with heterogeneous treatment effects (e.g., Su et al., 2009; Lipkovich et al., 2011; Loh et al., 2015), of which causal trees (Athey & Imbens, 2016) are arguably the most popular. [^1] Citation count was used as a proxy for popularity. In comparing the citation count of the different papers, as of 2025, the original causal tree paper by Athey & Imbens (2016) has over 2,000 citations, while the alternative methods have anywhere from 100-500 citations. Similar to decision trees (Breiman et al., 1984), causal trees partition the underlying covariate space into groups by finding clusters of individuals who have similar treatment effects. Moreover, unlike linear regression approaches, causal trees output a readily-interpretable partition of subgroups while also accounting for potential high-order interactions.
@@
In the following paper, we propose causal distillation trees (CDT) to stably estimate interpretable subgroups. CDT allows researchers to leverage the power of black-box meta learners, while preserving the interpretability of a simple tree output. CDT is a two-stage learner. The first stage estimates a heterogeneous treatment effect model using a flexible metalearner to generate a predicted heterogeneous treatment effect for each unit. The second stage then distills the information from the predicted heterogeneous treatment effect through a decision tree to construct interpretable subgroups. Intuitively, the first step serves as a de-noising step to help mitigate the instabilities that decision trees traditionally suffer from. The second step of ‘distillation’ draws upon existing ideas in machine learning (e.g., Hinton, 2015; Menon et al., 2021; Dao et al., 2021). However, unlike typical machine learning settings, where the primary goal is to improve the second stage learner’s predictive ability, we leverage distillation to improve our ability to estimate interpretable subgroups using the second stage decision tree learner.
@@ 2 Setup and Notation
Throughout the paper, we will use the potential outcomes framework, where $Y_{i}(1)$, $Y_{i}(0)\in\mathbbm{R}$ denote the potential outcome under treatment and control, respectively. Let $Z_{i}\in\{0,\;1\}$ denote the treatment assignment indicator. We assume consistency of treatment assignment across the study and no interference (i.e., the stable unit treatment value assumption). Let $Y_{i}:=Y_{i}(1)Z_{i}+Y_{i}(0)(1-Z_{i})$ be the observed outcomes. Finally, we assume that researchers have access to a set of pre-treatment covariates $X_{i}\in\mathbbm{R}^{p}$ for each unit in their study, and $\{Y_{i}(1),\;Y_{i}(0),\;X_{i},\;Z_{i}\}$ are drawn independently and identically distributed from an arbitrary joint distribution for i=1,,ni=1,...,n.
@@
$\tau_{\textsc{ATE}}=\mathbb{E}\left\{Y_{i}(1)-Y_{i}(0)\right\}.$
@@
where $\Omega$ represents a class of partitions in the covariate space of $X$. Under squared loss, the optimal partition corresponds to a subset in the covariate space of $X$ that maximizes the difference in treatment effect heterogeneity across subgroups. In other words, within a partitioned group, the treatment effect heterogeneity is minimized, thereby maximizing the variance in treatment effects across groups. In practice, researchers can choose alternative loss functions that correspond to what are substantively meaningful partitions (or subgroups) in the covariate space. However, the theoretical guarantees we derive utilize the squared loss, and generalizing the results for other loss functions is an open avenue of future research.
@@ 3.1 Overview
We propose a method, causal distillation trees (CDT), which stably estimates interpretable subgroups. CDT is a two-stage learner (Figure 1 and Algorithm 1). The first stage learner, referred to as the teacher model, learns an informative model for the heterogeneous treatment effect using the observed data $\{Y,\;Z,\;X\}$. Using this teacher model, we generate a prediction $\hat{\tau}_{i}^{d}$ for the heterogeneous treatment effect for all units. [^2] With slight abuse of notation, $\hat{\tau}_{i}^{d}$ corresponds to the predicted conditional average treatment effect for unit $i$ (i.e., for unit $i$, with covariates $X_{i}=x$, $\hat{\tau}_{i}^{d}:=\hat{\tau}(x)$) In the second stage, we distill the predicted $\hat{\tau}_{i}^{d}$’s into interpretable subgroups by fitting a decision tree, called the student model, to predict the $\hat{\tau}_{i}^{d}$’s from the covariates $X$. The estimated decision rules from the decision tree define a discrete partition, which naturally maps to the collection of subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\;\ldots,\;\hat{\mathcal{G}}_{G}(X)\right\}$ (see Figure A1). Finally, using the estimated partitions, we estimate the subgroup ATE for each subgroup.
Figure 1: Overview of Causal Distillation Trees (CDT). CDT leverages a two-stage procedure, which first fits a teacher model to estimate heterogeneous treatment effects and then fits a student model (e.g., a decision tree) to distill the estimated heterogeneous treatment effects and produce interpretable subgroups. The two-stage learner is learned using the training data (blue-gray boxes). Using the estimated subgroups, the subgroup ATEs are honestly estimated with a held-out estimation set (yellow box).
Informally, the first stage learner smooths the heterogeneous treatment effects by projecting $\tau_{i}$ into the basis space of the covariates $X$. By using the projected version of the heterogeneous treatment effect instead of $\tau_{i}$, the second stage learner will be able to more stably estimate the subgroups, as the teacher model will have de-noised the outcomes.
@@ 3.2 Consistency of subgroup estimation: Single rule setting
Let $F_{X}$ be the cumulative distribution function of the covariate $X$. Define the points $s_{l}$ and $s_{u}$ such that $0<F_{X}(s_{l})<F_{X}(s_{u})<1$. Let $s$ be an optimal split as defined in Definition 3.1. Assume that (i) $F_{X}$ is absolutely continuous with a density $dF_{X}$ that is bounded away from zero and is continuous in a neighborhood of $s$, and (ii) $s$ is unique.
@@
Under validity of the teacher model, the population-level optimal subgroups remain unchanged after distilling the individual-level treatment effects. In other words, the optimal population splits from the predicted heterogeneous treatment effects $\hat{\tau}_{i}^{d}$ are equivalent to the optimal split with the true individual-level treatment effects $\tau_{i}$. A sufficient, but not necessary, condition for a valid teacher model is consistency in recovering the conditional average treatment effects, which has been proven to be satisfied by many existing metalearners (see, e.g., Nie & Wager, 2021; Kennedy et al., 2022 for more discussion). Assumption (3.2) can also be relaxed to consider approximately valid teacher models. We provide additional discussion in Appendix A.2.
@@
[…]
  • For $\epsilon>0$, there exists an $1<\alpha<2$ where: $\inf_{|s^{\prime}-s|<\epsilon,s\neq s^{\prime}}\frac{|H(s)-H(s^{\prime})|}{|s-s^{\prime}|^{\alpha}}>0$.
@@
Proposition 3.1 formalizes that the rate of convergence depends directly on how smooth the underlying data generating process is. The rate of convergence in the estimated subgroups inherits a characteristic highlighted in Escanciano (2020), which bridged the gap in convergence rate of the estimated split in a decision tree $\hat{s}_{n}$ between a fully continuous setting (considered in Banerjee & McKeague, 2007 and Buhlmann & Yu, 2002) and discontinuous settings (considered in Chan, 1993; Kosorok, 2008).
@@
In addition to consistently recovering the subgroups, CDT can further improve the stability of the split estimation by de-noising the individual-level treatment effect and increasing the signal strength. To illustrate, in the following example, we consider a setting in which researchers have access to the true individual-level treatment effect $\tau_{i}$ and estimate a decision tree using the covariates $X_{i}$. In practice, this is infeasible, as researchers can never observe $\tau_{i}$, but this comparison serves as a helpful benchmark to compare the potential improvements from distillation.
@@ 3.3 General setting with multiple rules
For each subgroup index $g\in\{1,\ldots,G\}$ and rule index $r\in\{1,\ldots,r_{g}\}$, without loss of generality, write the subgroup defined by the r1r-1 previously selected binary decision rules as $\mathcal{G}_{g}^{(r-1)}(X)=\prod_{k=1}^{r-1}\mathbbm{1}\left\{X^{(j_{k})}\lesseqgtr s^{(j_{k})}\right\}$. We assume that there is separability, conditional on the previously selected binary decision rules:
@@
Assumption (3.5) ensures that the underlying subgroups have an optimal substructure so that a greedy second-stage decision tree learner is valid. If Assumption (3.5) is not met, standard greedy tree algorithms like CART can fail to recover the optimal partitions even with infinite amounts of data (see Tan et al., 2024 for a formal discussion). Recent work has introduced alternative algorithms to construct optimal trees under weaker assumptions, but at a much higher computational cost, and are often restricted to specific settings such as binary features, with limited out-of-sample performance gains relative to greedy trees (e.g., Hu et al., 2019; van der Linden et al., 2024). Given the popularity of greedy search approaches and the computational burden of many optimal tree construction approaches, we focus our discussion on settings in which greedy search algorithms can feasibly construct optimal subgroups.
@@
[…]
@@
The results of Theorem 3.1 follow from first noting that the estimated splits will converge in probability to the optimal split (see Lemma 3.1), and then showing that the within-sample loss will converge in probability to the true population-level loss (see Appendix B). In particular, the rate of convergence of $\hat{\mathcal{G}}_{g}(X_{i})$ to $\mathcal{G}_{g}(X_{i})$ is dependent on (1) the degree of separability in the losses between relevant and irrelevant rules (i.e., $\delta$), and (2) the underlying convergence rate of the sample splits to the true optimal splits (i.e., the rate at which $\left\lvert\hat{s}_{n}^{(k)}-s^{(k)}\right\rvert\to 0$). An immediate implication of Theorem 3.1 is that not only will the tree consistently split at the correct cutoff points, but it will also correctly select the important features that should be used in the splitting criteria.
An immediate implication of Theorem 3.1 is that for a fixed $G$, as $n\to\infty$, the true positive rate of the subgroups obtained under CDT will converge to 1—i.e.,
$\lim_{n\to\infty}\frac{1}{\lvert G\rvert}\sum_{g=1}^{G}\mathbbm{1}\{\hat{\mathcal{G}}_{g}(X)=\mathcal{G}_{g}(X)\}\to 1,$
and similarly, the false positive rate will converge to 0.
@@
Figure 2: Comparing the performance of causal trees versus CDT (using causal forests as the teacher model), measured via (A) number of estimated subgroups as well as (B) number of true positive and (C) number of false positive features used in the estimated subgroups. The oracle number of subgroups, true positives, and false positives are shown as dashed gray lines. Results are averaged across 100 simulation replicates with ribbons denoting $\pm 1SE$.
@@ 3.4 Estimating subgroup average treatment effects
Under Assumptions (3.1)-(3.7), the subgroup difference-in-means estimator will be a consistent estimator for the sample subgroup ATE. Furthermore, assume the proportion of units in each subgroup converges to a proportion $\pi^{*}_{g}\in(0,1)$ (i.e., $n_{g}/n\to\pi^{*}_{g}$ ad $n\to\infty$, where $0<\pi^{*}_{g}<1$ for all $g\in\{1,...,G\}$ and $\sum_{g=1}^{G}\pi^{*}_{g}=1$). Then, the asymptotic variance of $\hat{\tau}^{(g)}$ is
$\text{asyvar}(\hat{\tau}^{(g)})=\tilde{\beta}_{g}(1)\text{var}_{g}\{Y_{i}(1)\}+\tilde{\beta}_{g}(0)\text{var}_{g}\{Y_{i}(0)\},$
[…]
From Theorem 3.2, we can also construct a variance estimator using the sample analogs for $\text{var}_{g}(Y_{i}(z))$ for $z\in\{0,1\}$:
@@
where $\hat{\beta}_{g}(z)=\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\right)^{-1}$. Miratrix et al. (2013) showed that the gap between $\hat{\beta}_{g}(z)$ and $\tilde{\beta}_{g}(z)$ can be upper bounded as a function of $1/n$. Asymptotically, this will be equivalent to the robust Huber–White variance estimator.
We can apply standard nonparametric tests of treatment effect heterogeneity (see Appendix A). If researchers are interested in evaluating whether the largest treatment effect across the subgroups is statistically significantly different than the other treatment effects, we caution that a different test is needed due to post-selection bias. See Wei et al. (2024) for more discussion on valid re-sampling procedures for inference in these settings.
If researchers are interested in the finite-population subgroup ATE, then we can show that the subgroup difference-in-means estimator will be an unbiased estimator for the finite-population subgroup ATE. However, the variance estimator introduced above will be a conservative estimator of the true variance in a finite-population setting. (See Appendix A.4 for more discussion.)